OpenStudy (anonymous):

what is a race car's acceleration if its velocity changes from 0 m/s to 40 m/s in 4 seconds.

OpenStudy (whpalmer4):

\[v_f - v_i = a t\]where \(v_i\) is initial velocity, \(v_f\) is final velocity, \(a\) is (constant) acceleration, and \(t\) is time

OpenStudy (anonymous):

so 10 m/s

OpenStudy (whpalmer4):

Not quite. Right number, wrong unit, thus wrong answer...

OpenStudy (whpalmer4):

As \(v_i=0\), \[v_f = at\]\[a = \frac{v_f}{t}\] \[a = \frac{ 40\frac{\text{ m}}{\text{ s}} }{4 \text{ s}} = 10 \frac{\text{m}}{\text{s}^2}\]

OpenStudy (whpalmer4):

acceleration always has dimensions of \[\frac{\text{[length]}}{\text{[time]}^2}\]

OpenStudy (anonymous):

Ahhh okay

OpenStudy (whpalmer4):

Always do your physics problems with the units. If the units don't behave properly, you've set the problem up wrong, or muffed a unit conversion, and your answer is undoubtedly wrong.

OpenStudy (anonymous):

Okay so just 10 m/s^2

OpenStudy (whpalmer4):

Yes, that is correct. Coincidentally, that's about the same acceleration as gravity on the earth's surface (9.81 m/s^2), so you'll sometimes see that referred to as "1 g"

OpenStudy (anonymous):

Thank you